OpenStudy (anonymous):

Judy works at a local bowling alley. She currently makes $300 per week after taxes plus $60 extra cash every other week for cleaning the lanes. Judy is trying to save up for a car. She really wants a Bug. She knows that she can buy a used Bug in good condition for between $6,600 and $9,900. The equation below represents this situation where x is the number of weeks worked and P is the price of the car: What is the minimum number of weeks that Judy needs to work to buy her Bug? 30 weeks 20 weeks 25 weeks 10 weeks

OpenStudy (anonymous):

guess we can average out the extra $60 every two weeks and make it $30 per week. then she makes a total of \(300+30=\$330\) per week

OpenStudy (anonymous):

so if she works \(x\) weeks she will make \(330x\) dollars

OpenStudy (anonymous):

and you get to solve \(6,600\leq 330x\leq 9,900\) for \(x\) divide across by \(330\) and be done in one step